### v1

# Study of transport properties of a hot and dense QCD matter using a novel approximation method

Abstract
We have studied the charge and heat transport properties of a hot and dense QCD matter using a novel approximation method within the quasiparticle model. Utilizing a novel collision integral for both the relaxation time approximation (RTA) and the Bhatnagar-Gross-Krook (BGK) models, we have solved the relativistic Boltzmann transport equation to estimate the electrical conductivity and the thermal conductivity. We have investigated the temperature dependence of these transport coefficients. We have also provided a comparison between our findings and those of the standard RTA and BGK models. Further, we have explored the temperature dependence of the thermal diffusion constant and the Lorenz number using the novel approaches of the aforesaid models.
1 Introduction
A long time goal of the physics of ultrarelativistic heavy ion collisions is the observation of new states of matter predicted by the quantum chromodynamics (QCD). The experimental evidences of the relativistic heavy ion collider (RHIC) [1] at BNL and the large hadron collider (LHC) [2] at CERN indicate a new state of strongly interacting matter known as quark-gluon plasma (QGP). Following the discovery of QGP, various properties of the strongly coupled QCD matter became a topic of high interest. Different research groups have studied different features of QGP in recent years, such as the thermodynamic properties [3, 4], the relativistic hydrodynamics [5, 6, 7], the heavy quark diffusion in QCD matter [8, 9, 10], the memory effect on heavy quark dynamics [11, 12], the dissipative spin dynamics in relativistic matter [13, 14] etc. Nowadays, there is much interest in the phase diagram region at finite temperature and baryon chemical potential [15, 16]. Many observations have also been made recently to investigate the impact of a strong magnetic field on the characteristics of hot QCD matter, such as the chiral magnetic effect [17, 18], the chiral vortical effect [19], the axial Hall current [20], the refractive indices and decay constant [21, 22], the photon and dilepton productions from QGP [23, 24, 25], the thermodynamic and magnetic properties [26, 27, 28, 29, 30, 31], the axial magnetic effect [32, 33], the thermoelectric responses of hot QCD medium in a time varying magnetic field [34, 35], the magnetic and inverse magnetic catalysis [36, 37], the nonlinear electromagnetic current [38, 39] and the magnetohydrodynamics [40, 41].
It is crucial to investigate different transport coefficients since they shed light on the genesis of the hot QCD medium. Among these transport coefficients, the thermal conductivity () and the electrical conductivity () characterize the movement of heat and charges in the medium, respectively. The hydrodynamic development of the strongly interacting matter at finite baryon density also depends on these two conductivities [42, 43]. Therefore, significant effects on the observables at heavy ion collisions might result from changes in the charge and heat transport coefficients. There are several methods for determining these conductivities, such as the relativistic Boltzmann transport equation [44, 45, 46], the Chapman-Enskog approximation [47, 48], the lattice simulation [49, 50, 51], the correlator technique using Green-Kubo formula [52, 53, 54] etc. Previous research has examined the impact of magnetic field on the electrical conductivity using a variety of methods, including the diluted instanton-liquid model [52], the quenched SU(2) lattice gauge theory [55], the effective fugacity technique [56], the diagrammatic method utilizing the real-time formalism [57] etc. Studies on the electrical and thermal conductivities of hot QCD matter in the presence of magnetic field within the kinetic theory approach have been carried out in references [58, 59, 60, 61].
The application of theoretical methodologies intended for systems of this type has been made viable by the strongly coupled fluid behavior of QGP. A simplified version of the kinetic equation with a simplified form of the collision term is frequently used because of various technical limitations. It is referred as the relaxation time approximation (RTA) [62]. In this model, the mean-free approach reduced the complexity of the collision term. The development in the nonrelativistic case has already been achieved by Bhatnagar-Gross-Krook (BGK) model [63] and Welander model [64]. This approximation is utilized in many domains of physics and is recently used to investigate the hydrodynamization of the matter generated by the ultrarelativistic collisions of heavy ions. The relaxation time approximation has become essential for exploring the domain and microscopic basis of the relevance of hydrodynamics with relativistic effects. Although widely used, the approximation presented in ref. [62] has some fundamental challenges. It conflicts with some macroscopic and microscopic conservation laws. Recently, a paper [65] provided a novel relaxation time approximation that fixes such fundamental issues and maintains the core features of the linearized Boltzmann collision operator. In light of the momentum-dependent relaxation time, the novel relaxation time approximation is significant as it maintains the fundamental collision features associated with the microscopic conservation laws. They showed how this adjustment impacts the particle diffusion and the bulk viscosity. In a recent study [66], the transport coefficients related to charge and heat for a rotating QGP medium have been derived within the RTA model utilizing the aforementioned collision integral.
In this study, we estimate the response functions of charge flow and heat flow, such as the electrical conductivity () and the thermal conductivity (), to investigate the charge and the thermal transport properties of the hot and dense QCD matter. These transport coefficients have been computed using a modified collision integral in the relativistic Boltzmann transport equation within the kinetic theory approach. We also explore the thermal diffusion constant and the Lorenz number as the applications of charge and heat transport properties in an analogous environment. The thermal diffusion constant shows the overall rate of heat transport in the medium, while the Lorenz number illustrates the connection between the thermal and electrical conductivities. We use the quasiparticle description of partons, where the rest masses are replaced by the masses generated in the medium. The quark-gluon plasma is viewed as a medium made up of thermally massive quasiparticles in this framework [67, 68].
The rest of the paper has been organized as follows: Using the modified collision integral, the charge and heat transport coefficients from the relativistic Boltzmann transport equation are calculated in section 2. In section 3, we have discussed the results of the aforesaid transport coefficients using the quasiparticle description of partons. Section 4 describes various applications and their results obtained from these coefficients, such as the Lorenz number and the thermal diffusion constant. In section 5, the results are summarized.
Notations and conventions: The covariant derivatives and denote and respectively. The fluid four-velocity in the rest frame () is normalized to unity. In this work, the subscript stands for the flavor index, with =u, d, s. The symbols , and () stand for electric charge, degeneracy factor and the infinitesimal change in the distribution function for the quark (antiquark) of the flavor, respectively. Here, represents the degeneracy factor for the gluonic degree of freedom, and denotes the degeneracy factor for the quark (antiquark) degree of freedom with flavor , where is the number of colours.
2 Charge and heat transport properties of a hot and dense QCD medium
For a single particle distribution function, the relativistic Boltzmann transport equation (RBTE) is written as
where F is the force field acting on the particles in the medium and is the mass of the particle. The term on the right-hand side results from particle collisions within the medium. This term’s zero value denotes the Vlasov equation, which describes a collisionless system. The intricate character of the collision term presents a number of challenges even for the simplest solution of the above eq. (1). The term represents the instantaneous change in the distribution function due to collisions between particles. In order to arrive at a comprehensive solution, we will first go over a few mathematical models that allow us to address the collision term in a suitable manner.
When using the relaxation time approximation (RTA), the collision term is substituted with a relaxation term of the following form,
where we can write and is the relaxation time for collisions, i.e. the time required to bring the perturbed system back to its equilibrium state. We can suppose the quark distribution function is near equilibrium with a slight divergence from equilibrium in order to permit linearization of the Boltzmann equation. Here, the term acts as a damping frequency. The initial issue with this model seems to be the immediate conservation of charge. To remove the crunch, the Bhatnagar-Gross-Krook (BGK) model presents a new type of collision kernel, denoted as
where represents the equilibrium density and is known as the perturbed density or the fluctuating density. The perturbed density is calculated as which gives zero value after integrating over momentum, thus indicating instantaneous conservation of the particle number during collisions in the system. Such fundamental problems are resolved and the essential characteristics of the linearized Boltzmann collision operator are preserved via a novel relaxation time approximation that was recently presented in a research [65].
Here, we are considering the general relativistic covariant form of the relativistic Boltzmann transport equation for an isotropic medium of quarks, antiquarks and gluons, which is given as
where is the electromagnetic field strength tensor. In order to see the effects of electric field, we take only and and vice versa, components of the electromagnetic field strength tensor, i.e. and in our calculation, thus the RBTE (4) takes the following form,
The infinitesimal change in quark distribution function is denoted as and indicates the equilibrium distribution function in the medium for flavor.
2.1 Electrical conductivity in the novel RTA and the novel BGK models
The electric current density of the medium must be determined in order to see how the electric field affects the charge transport phenomena. An external electric field can cause an infinitesimal disturbance to a thermal QCD medium that contains quarks, antiquarks, and gluons of distinct flavors. Due to this electric field, there is an induced current density in the medium, which is given by
where represents the spatial part of the electric current density vector. Ohm’s law states that the spatial component of the electric current density is directly proportional to the electric field with a proportionality factor, known as the electrical conductivity (). Thus, we have
By comparing eq. (6) and eq. (7), one can obtain the expression of the electrical conductivity.
By solving the relativistic Boltzmann transport equation, the infinitesimal changes in the quark and antiquark distribution functions can be obtained. The relativistic Boltzmann transport equation in the novel-RTA model can be written as
where the collision integral is given [65] by
Here,
and we define the momentum integrals relative to the local equilibrium distribution function as follows,
where is the relaxation time and denotes the collision frequency of the medium. The equilibrium distribution functions for flavor of quark and antiquark are given by
respectively, where , , and represents the chemical potential of flavor of quark. Getting inspired by [69, 70, 71], we have used the following expression of the relaxation time,
where is an arbitrary constant which controls the energy dependence of the relaxation time and the expression for the is given [72] by
Here, is the QCD running coupling constant, which is a function of both temperature and chemical potential and it has the following [73] form,
where GeV, for electrically charged particles (quarks and antiquarks) and for gluons. Using the expression of in eq. (8), we have
Now, solving the above equation one can extract the expression of (for a detailed calculation see appendix A) as
where
Similarly for antiquarks,
where
Finally, replacing the values of and in eq. (6) and comparing it with eq. (7), we get the final expression of the electrical conductivity in the novel RTA model as
In the novel-BGK model, the relativistic Boltzmann transport equation can be expressed as
Now, neglecting higher order terms , we get the solution of this equation as
where
The infinitesimal change in the quark distribution function is obtained (details are given in appendix B) as
Similarly, is obtained as
Now, replacing the values of and in eq. (6) and comparing it with eq. (7), we get the final expression of the electrical conductivity in the novel BGK model as
2.2 Thermal conductivity in the novel RTA and the novel BGK models
Heat flow results from the deviation of the system from its equilibrium condition caused by the presence of a temperature gradient. Through a proportionality constant, i.e. the thermal conductivity, it is directly proportional to the temperature gradient. The calculation of the thermal conductivity is necessary in examining the heat transmission in a medium. In four-vector notation, the heat flow is written as
In the above equation, the projection operator () and the enthalpy per particle () are respectively defined as
where , and represent the energy density, the pressure and the particle number density, respectively, which are respectively given by the following equations,
Here, the particle flow four-vector and the energy-momentum tensor are respectively defined as
The spatial part of the heat flow four-vector is given by
where we use the fact that in the rest frame of the fluid, the heat flow four-vector and fluid four-vector are perpendicular to each other, i.e. and the enthalpy per particle for the flavor is given by
where
In the Navier-Stokes equation, the heat flow is related to the gradients of temperature and pressure [74] as
Here, represents the thermal conductivity. One can find the expression of the thermal conductivity by comparing eq. (39) and eq. (43).
In order to obtain the expression of thermal conductivity, we start with the relativistic Boltzmann transport equation, which can be rewritten as
With the help of eq. (9), the above equation can be written in the novel RTA model as
Now, solving this equation one can extract the expression of (for a detailed calculation see appendix C) as
Similarly for antiquarks,
Finally, replacing the values of and in eq. (39) and comparing it with eq. (43), we get the final expression of the thermal conductivity in the novel RTA model as
In the novel-BGK model, the relativistic Boltzmann transport equation takes the following form,
Now, neglecting higher order terms , we get the solution of this equation as
where
The infinitesimal change in the quark distribution function is obtained (for a detailed calculation see appendix D) as
Similarly, is obtained as
Finally, replacing the values of and in eq. (39) and comparing it with eq. (43), we get the final expression of the thermal conductivity in the novel BGK model as
3 Results and discussions
The interaction of the partons with the thermal medium gives rise to a certain mass known as the thermal mass or the quasiparticle mass. Within the framework of the quasiparticle model, we may see QGP as a system of massive noninteracting quasiparticles. The quasiparticle quark-gluon plasma (qQGP) model is a commonly employed framework to explain the collective properties of the QGP medium. The quasiparticle masses of particles in a dense thermal medium rely on both temperature and chemical potential. For a dense QCD medium, the thermal masses (squared) of quark and gluon up to one-loop are given [67, 68] by
respectively, with , where the expression of is given in eq. (18). It can be seen from eq. (18) that, depends on both temperature and chemical potential. In this work, we have used GeV. The accessible energy scales for a thermal QCD medium are related to the temperature, quark mass and chemical potential. The QGP phase, which is made up of light quarks, light antiquarks and gluons, has a temperature high enough to reach the maximum energy scale. The hard thermal loop (HTL) perturbation theory is one of the effective theories that is applied to tackle the divergences that arise in the computation of the QCD thermodynamic observables, the transport coefficients and the amplitudes in the high temperature regime.
In this section, the charge and heat transport coefficients are investigated with the thermal masses as given in equations (55) and (56). In the novel RTA and the novel BGK models, quantities and appear in the denominators of the electrical conductivity and the thermal conductivity expressions. Thus the knowledge on the variations of these quantities with temperature and momentum is important to understand the variations of the conductivities. In figures 1 and 2, we have plotted and for momentum-dependent relaxation time () as well as for momentum-independent relaxation time (), respectively. We have found decreasing trends of and with the increase of temperature (figures 1(a) and 2(a)) and their increasing trends with the increase of momentum (figures 1(b) and 2(b)). In both the cases, it is observed that the momentum dependence of the relaxation time enhances both and . Thus, through the effects of the momentum-dependent relaxation time on and , it is possible to comprehend how the momentum-dependent relaxation time would modulate the transport coefficients.
Figures 3(a) and 3(b) show the variations of the electrical conductivity () with the temperature in the novel RTA and the novel BGK models at finite chemical potential, respectively. More precisely, we have compared our results on the electrical conductivity with those of the standard RTA and BGK models. We have also compared the result for with that for . The relaxation time appears to be momentum-dependent (eq. (16)) in the first case, whereas the relaxation time is momentum-independent in the second case. We have observed that, regardless of the relaxation type ( or ), there is a noticeable reduction in the magnitudes of in the novel RTA and the novel BGK models as compared to those in the RTA and the BGK models, respectively. This reduction in the electrical conductivity in the novel RTA and the novel BGK models is mainly attributed to the quantities and appearing in the denominators of equations (24) and (30).
T (0.16 GeV) | ||||||||
0.50 | 0.51 | 0.64 | 0.78 | 0.64 | 0.78 | 1.26 | 1.51 | |
0.40 | 0.45 | 0.45 | 0.57 | 0.45 | 0.57 | 1.11 | 1.27 |
T (1 GeV) | ||||||||
0.52 | 0.54 | 0.54 | 0.54 | 0.68 | 0.83 | 1.55 | 1.30 | |
0.41 | 0.46 | 0.46 | 0.46 | 0.47 | 0.60 | 1.29 | 1.26 |
In figures 4(a) and 4(b), the thermal conductivity () has been plotted as a function of the temperature in the novel RTA and the novel BGK models at finite chemical potential, respectively. These figures show the comparison of our results on the thermal conductivity with those of the standard RTA and BGK models. We have found that, irrespective of the relaxation type ( or ), there is a discernible decrease in the magnitudes of in the novel RTA and the novel BGK models as compared to those in the RTA and the BGK models, respectively. The quantities and appearing in the denominators of equations (48) and (54) mainly contribute in the decrease of the thermal conductivity in the novel RTA and the novel BGK models.
The electrical and thermal conductivities are found to be significantly impacted by the adjustments we made to the RTA model as well as the BGK model when an energy-dependent relaxation time () is used. Tables 1 and 2 show how the updated collision integral affects the charge and heat transfer in general in the low temperature and the high temperature regimes, respectively. Based on the electrical conductivity related ratios in tables 1 and 2, we can conclude that, for the novel RTA and the novel BGK models, the drop in the magnitude of becomes larger when considering the momentum dependent relaxation time as compared to the standard RTA and BGK models. We may infer from the thermal conductivity related ratios in tables 1 and 2 that, as compared to the standard RTA and BGK models, the reduction in the magnitude of for the novel RTA and the novel BGK models increases when the momentum dependent relaxation time is taken into account. Thus the momentum dependence of the relaxation time decelerates the flow of charge as well as the flow of heat in the medium.
4 Applications
The purpose of this section is to study some applications of electrical and thermal conductivities for the hot and dense QCD medium. In subsection 4.1, we analyze the heat transfer in the medium through the thermal diffusion constant in the novel RTA and the novel BGK models. The relative behavior of the thermal conductivity and the electrical conductivity is explored in subsection 4.2 using the Lorenz number in the aforesaid models.
4.1 Thermal diffusion constant
Instead of being measured directly, the thermal conductivity in common materials is usually determined from measurements of the thermal diffusion constant . This is related to the rate of heat transfer in the medium. The thermal diffusion constant in the first order relativistic viscous hydrodynamics is defined as
Here, represents the specific heat at constant pressure and it can be determined by using the following formula.
where and are given in equations (34) and (35), respectively. Increase of the thermal diffusion constant explains the faster heat transfer in the medium.
Figure 5 shows the variation of the thermal diffusion constant with the temperature at a fixed chemical potential in the novel RTA and the novel BGK models. It is evident from figures 5(a) and 5(b) that the magnitude of gets decreased regardless of the relaxation type ( or ) for both RTA and BGK models with the modified integral term. When we compare it with the standard RTA and BGK models, we have found that the momentum-dependent relaxation time () leads to a considerable shift in the thermal diffusion constant. We can conclude from the thermal diffusion constant related ratios in tables 1 and 2 that, for the novel RTA and the novel BGK models, the decline in the magnitude of becomes larger when considering the momentum-dependent relaxation time as compared to the standard RTA and BGK models.
4.2 Lorenz number
The Lorenz number () is the proportionality factor in the Wiedemann-Franz law, which states that the ratio of the heat transport coefficient () to the charge transport coefficient () is directly proportional to the temperature, i.e.,
This law describes how the phenomena of heat transfer and charge transfer behave in relation to one another in a medium.
Figure 6 shows the variation of the Lorenz number with temperature at a fixed chemical potential. It is apparent from figures 6(a) and 6(b) that the Lorenz number exhibits a notable increase in magnitude in case of the novel RTA and the novel BGK models, regardless of the relaxation type ( or ). A significant change in the Lorenz number is also observed as a result of the momentum dependence of the relaxation time () for both RTA and BGK models. From the Lorenz number related ratios in tables 1 and 2, it is observed that, the increase in the magnitude of for the novel RTA and the novel BGK models as compared to the normal RTA and BGK models gets smaller when the momentum-dependent relaxation time is used.
5 Summary
In this work, we studied the transport properties of a hot and dense QCD matter in the novel RTA and the novel BGK models. This study could help to get an understanding of the charge and heat transport phenomena in a strongly coupled thermal QCD medium with a finite chemical potential background through the novel approaches of the aforesaid models. Using the kinetic theory technique, we solved the relativistic Boltzmann transport equation with a modified collision integral for estimating the thermal and electrical conductivities of the aforementioned medium. We examined the effects of the modified collision integral on the charge and heat transport in a thermal QCD medium, in contrast to the widely used collision terms of the RTA and the BGK models. In this research, we used the quasiparticle description of partons, where both the temperature and the chemical potential affect the constituent masses in the medium. Our modifications to the collision integral for the RTA and the BGK models were shown to result in a considerable drop in the electrical and thermal conductivities. Additionally, we found that these transport coefficients exhibit a discernible shift when we take into account a momentum-dependent relaxation time (). In the same regime, the thermal diffusion constant and the Lorenz number were also investigated. We observed that the novel collision integral leads to an increase in the Lorenz number and a decrease in the thermal diffusion constant of the medium.
6 Acknowledgments
One of us (S. R.) is thankful to the Indian Institute of Technology Bombay for the Institute postdoctoral fellowship and S. D. acknowledges the SERB Power Fellowship, SPF/2022/000014 for the support on this work.
Appendix A Derivation of equation (20)
Using the result of ref. [65], one can write the relativistic Boltzmann transport equation as
The second term in the right-hand side can be reduced to
Similarly, the 3rd and 4th terms can be respectively reduced to
Now, with the help of equations (A.61), (A.62) and (A.63), eq. (A.60) takes the following form,
where,
Using the following partial derivatives,
we solve eq. (A.64) to get the expression of as given in eq. (20).
Appendix B Derivation of equation (28)
In the BGK model, the relativistic Boltzmann transport equation is written as
For a proper treatment of the right-hand side of the above equation, we are considering an infinitesimal perturbation to the equilibrium distribution function, i.e. with and after linearizing it, we get
Now, incorporating the result of ref. [65] in the novel BGK model, we get
Using eq. (B), eq. (B.69) and the partial derivatives as given in eq. (A.66), eq. (B.67) takes the following form,
After some mathematical calculation, we get
where is related to the infinitesimal change of the distribution function () by the relation given in eq. (26). After putting the value of in eq. (26), we get as given in eq. (28).
Appendix C Derivation of equation (46)
Appendix D Derivation of equation (52)
In the BGK model, the relativistic Boltzmann transport equation is written as
With the help of eq. (B), eq. (B.69) and the partial derivatives as given in eq. (C.73), we can express the above equation in the novel BGK model as
After some mathematical calculation, we get
where is related to the infinitesimal change of the distribution function () by the relation given in eq. (50). After putting the value of in eq. (50), the value of , i.e. eq. (52) is obtained.